Two firms, A and B, both produce brushes. The price of brushes is $1.50 each. Firm A has total fixed costs of $450,600 and variable costs of 54 cents per brush. Firm B has total fixed costs of $260,000 and variable costs of 72 cents per brush. The corporate tax rate is 30%. If the economy is strong, each firm will sell 1,506,000 brushes. If the economy enters a recession, each firm will sell 977,000 brushes.

Calculate Firm A's degree of operating leverage.
Answer #1

For Firm A

Let the probability of a good economy and bad economy be equal,

Hence, number of brushes sold = ΣProbability of Economy * Number of brushes sold = 0.50*1506000 + 0.50*977000 = 1241500

Total Sales = Number of brushes sold * Price of Brush = 1241500 * 1.50 = $1862250

Fixed Cost = $450600

Variable Cost = Number of brush sold * Variable cost/unit = 1241500 * 0.54 = $670410

Degree of Operating Leverage = (Sales - Variable Costs)/(Sales - Variable Costs - Fixed Costs)

= (1862250 - 670410) / (1862250 - 670410 - 450600) = 1.608